OpenStudy (anonymous):

18/20=k/110 plzzzz help me

OpenStudy (tyteen4a03):

Hint: Remember our trustworthy cross-multiplication: a/b = c/d can be written as ad = bc.

OpenStudy (anonymous):

Please see attached file for an explanation as to how to solve this problem.